calculate the change in volum of gases when 1 mole pf PH3 decompsed at STP according to given reaction.
PH3(g)---1/4P4(g)+3/2H2 (g)
how to slove this?

Grade:12

1 Answers

Ramreddy IIT Patna
9 years ago
Volume of gas is proportional to number of moles.
products volume= (1/4 + 3/2)
= 7/4 mL (lets take volume in mL)
reactans volume= 1 mL
Change in volume= 7/4-1 =3/4 mL
….….….….….….….….….….….….….….….….….….….….….…....................................